OpenStudy (anonymous):

log 10 x - log 10 (x-4) =1

OpenStudy (anonymous):

is it log of x to the base 10??

OpenStudy (anonymous):

x=40/9

OpenStudy (anonymous):

log10x-log10(x-4)=log(10x/10(x-4))=log(x/x-4)=1 hence, x/x-4=10 solving this u will get the answer as x=40/9